On Tuesday a hotel scheduled 9 workers, each of which worked 8 hours cleaning guest rooms and had a 30 minute break. On that day, 153 guest rooms were cleaned. What were the minutes per room achieved by these workers on Tuesday?
Select one:
a. 26 minutes
b. 22 minutes
c. 27 minutes
d. 17 minutes
Clear my choice


11)
Answer #1
Total number of workers (A) - given 9
Schedule hours per worker (B) - given 8 hrs
Total number of minutes available for all workers (C) = A*B*60mins 9*8*60 = 4320 mins
Break of 30 mins for each workers (D) = 9 workers * 30 mins 270 mins
Total number of minutes worked by all workers (E) = D-C 4320-270 = 4050 mins
Number of guest rooms cleaned (F) - given 153 rooms
Minutes per room (G) = E/F 4050 mins/153 rooms = 26.47 minutes, rounded to 26 minutes
Therefore, option A 26 minutes is the correct answer as calculated above
Other options are incorrect as they are either excessive or lesser minutes